# Self-Adaptive Physics-Informed Quantum Machine Learning for Solving Differential Equations

### v2

@@ 4 Discussion
Initially, a highly nontrivial Riccati equation is solved, and a system of differential equations is investigated comparing the results to existing solutions. In addition, a second-order differential equation is analyzed. The importance of balancing the multi-objective loss function is presented using constant balancing weights. To further explore the generalized case, an alternative approach based on self-adaptive weights is presented incorporating the optimization of balancing weights. Lastly, a second-order 2D partial differential equation is investigated. Therefore, a new quantum circuit is designed that can incorporate multiple independent variables.
Overall, the modified quantum circuits with added entangling layers lead to more accurate results in fewer epochs while solving second-order ordinary and partial differential equations. This work showcases the promise of quantum machine learning in addressing a variety of differential equations, offering an innovative and efficient solution approach compared to classical methods. The adaptability of the proposed framework to different problem domains and its potential for further extensions, such as 3D partial differential equations and diverse applications e.g. in engineering or materials science, underscore the significance of advancing the field of quantum computing for solving real-world problems. Nevertheless, many questions remain regarding the capabilities of the presented work when applying to different problems such as nonlinear mechanics e.g. modeling finite-strain plasticity, solving the Navier-Stokes equation, or even coupled problems for fluid-structure interaction. A further question that still needs to be addressed is if a potential quantum advantage can be achieved. Thus, the presented work is an initial starting point for further applications of QML presenting promising results to approach these problems.
@@
The datasets used and/or analyzed during the current study are available from the corresponding author on reasonable request

### v3

@@ 4 Discussion
Initially, a highly nontrivial Riccati equation is solved, and a system of differential equations is investigated comparing the results to existing solutions. In addition, a second-order linear differential equation is analyzed. The importance of balancing the multi-objective loss function is presented using constant balancing weights. To further explore the generalized case, an alternative approach based on self-adaptive weights is presented incorporating the optimization of balancing weights. Furthermore, a complex example of the nonlinear Duffing equation is analyzed. Lastly, a second-order 2D partial differential equation is investigated. Therefore, a new quantum circuit is designed that can incorporate multiple independent variables.
Overall, the modified quantum circuits with added entangling layers lead to more accurate results in fewer epochs while solving second-order ordinary and partial differential equations. The enhancement in performance can be attributed to the fact that quantum circuits containing a larger number of entangling layers will result in a higher degree of expressibility. Consequently, the quantum states within the Hilbert space will be more thoroughly explored, leading to more precise outcomes. This work showcases the promise of quantum machine learning in addressing a variety of differential equations, offering an innovative and efficient solution approach compared to classical methods. The adaptability of the proposed framework to different problem domains and its potential for further extensions, such as 3D partial differential equations and diverse applications e.g. in engineering or materials science, underscore the significance of advancing the field of quantum computing for solving real-world problems. Nevertheless, many questions remain regarding the capabilities of the presented work when applied to different problems such as nonlinear mechanics e.g. modeling finite-strain plasticity, solving the Navier-Stokes equation, or even coupled problems for fluid-structure interaction. A further question that still needs to be addressed is whether a potential quantum advantage can be achieved. Thus, the presented work is an initial starting point for further applications of QML presenting promising results to approach these problems.
@@
The datasets used and/or analyzed during the current study are available from the corresponding author upon reasonable request.